OpenStudy (anonymous):

on monday afternoon kelly withdrew 5% of the balance of her bank account to loan her friend some money. on tuesday morning she deposited 100.00 into her bank account she now has 25.00 more than she had on monday morning how much was in her bank account monday morning

OpenStudy (anonymous):

This is really interesting susie. Who is this Kelly you speak of?

OpenStudy (anonymous):

its a question that my instructor gave me to figure out so i don't know her

OpenStudy (anonymous):

kelly had 1500 in her accouint on monday morning

OpenStudy (anonymous):

how did you get this answer

OpenStudy (anonymous):

let the sum kelly had on Monday morning = x she withdrew 5% = 5/100 * x = 5x/100 now sum left in her account is x - (5x/100) = 95x/100 on Tuesday morning after depositing 100 she had 25 more than what she had on Monday morning => 95x/100 + 100 = x + 25 => 95x/100 - x = 25 - 100 95x - 100x => ------------ = -75 100 => - 5x ---- = -75 100 => -5x = -75 * 100 => x = -7500/-5 so x = 1500

OpenStudy (anonymous):

if u r satisfied with the answer, pls click GOOD ANSWER button for me :)

OpenStudy (anonymous):

thank you very much i had to step away for a minute thats what took so long

OpenStudy (anonymous):

no problem..... Thanks!!